OpenStudy (anonymous):

Ok, one more...limit as x approaches 0 of (e^x+x)^1/x?

OpenStudy (amistre64):

[x]' = 1 ... so thats covered :)

OpenStudy (amistre64):

err... is the exponent ^(1/x)?

OpenStudy (anonymous):

(1+0)^înf = 1

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=limit+as+x+approaches+0+of+%28e^x%2Bx%29^1%2Fx goes to e^2 if this is right

OpenStudy (anonymous):

lim_(x->0) (e^x+x) = 1

OpenStudy (dumbcow):

e^(x+x)^(1/x) = e^(2x)*(1/x) = e^2

OpenStudy (anonymous):

yep, the e^2 is the right answer

OpenStudy (anonymous):

e^2

OpenStudy (anonymous):

Just slot some figures in...